OpenStudy (anonymous):

Solve for h. A = 1/2 bh A.) 2a - b = h B.) -1/2a - b = h C.) a - b / 2 = h D.) 2a / b = h

OpenStudy (anonymous):

A = 1/2 bh 2A = bh h = 2a/b ans = D.